# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v2

@@ 6 Lower Bound
In this section, we establish a lower bound for MNL mixture MDPs by presenting a novel reduction, which connects MNL mixture MDPs and the logistic bandit problem.
Consider the following logistic bandit problem (Faury et al., 2020; Abeille et al., 2021): at each round $k\in[K]$, the learner selects an action $x_{k}\in\mathcal{X}$ and receives a reward $r_{k}$ sampled from Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*}))^{-1}$, where $\theta^{*}\in\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$ is the unknown parameter. The learner aims to to minimize the regret:
${\operatorname*{Reg}}^{\mathtt{LogB}}(K)=\max_{x\in\mathcal{X}}\sum_{k=1}^{K}\mu(x^{\top}\theta^{*})-\sum_{k=1}^{K}\mu(x_{k}^{\top}\theta^{*}).$
Theorem 3 (Lower Bound).
@@
Corollary 1 can be proved by combining Theorem 3 and the $\Omega(d\sqrt{K\kappa^{*}})$ lower bound for logistic bandits with infinite arms by Abeille et al. (2021). To the best of our knowledge, a lower bound for logistic bandits with finite arms has not been established, which is beyond the scope of this work. This absence leaves the lower bound for MNL mixture MDPs with a finite action space open through this reduction. However, after the submission of our work to arXiv (Li et al., 2024a), a follow up work by Park et al. (2024) proposed a new reduction that bridges MNL mixture MDPs with linear mixture MDPs by approximating multinomial logit functions to linear functions, employing the mean value theorem. Leveraging this new reduction, they established an $\Omega(dH^{3/2}\sqrt{K})$ lower bound for MNL mixture MDPs with the finite action space. This achievement confirms that our result is indeed optimal with respect to the dependence on $d$ and $K$, only differing by a $\mathcal{O}(H^{1/2})$ factor.
Dependence on $H$.  By the discussion in Remark 7, we note that our result is optimal with respect to $d$ and $K$, but loosing by a $\mathcal{O}(H^{1/2})$ factor. We discuss the challenges in improving the dependence on $H$. Notably, MNL mixture MDPs can be viewed as a generalization of linear mixture MDPs (Ayoub et al., 2020; Zhou et al., 2021). The pioneering work by Ayoub et al. (2020) achieved a regret bound of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K})$ for linear mixture MDPs, which matches our results in Theorem 2, differing only on the lower order term. Later, Zhou et al. (2021) enhanced the dependence on $H$ and attained an optimal regret bound of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$. This was made possible by recognizing that the value function in linear mixture MDPs is linear, allowing for direct learning of the value function while incorporating variance information. In contrast, the value function for MNL mixture MDPs does not conform to a specific structure, posing a significant challenge in using the variance information of value functions. Thus, it remains open whether similar improvements on $H$ are attainable for MNL mixture MDPs.
@@ 7 Conclusion and Future Work
In this work, we study MNL mixture MDPs that employ multinomial logit function approximation to ensure valid probability distributions over the state space. We address both the computational and statistical challenges for this problem. Specifically, we first propose an algorithm based on the online Newton step that attains the $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode. Next, we propose an enhanced algorithm that leverages local information to improve the statistical efficiency. It not only maintains $\mathcal{O}(1)$ computational and storage costs but also achieves an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, nearly matching the result of linear function approximation from both computational and statistical perspectives. Finally, we establish the first lower bound for MNL mixture MDPs, justifying the optimality of our results in $d$ and $K$.
There are several interesting directions for future work. First, there still exists a gap between the upper and lower bounds and how to close this gap is an open problem. Besides, while this work focuses on stochastic rewards, extending this model to the non-stationary settings and studying the dynamic regret (Wei and Luo, 2021; Zhao et al., 2022; Li et al., 2023) is an another important direction.
@@ Acknowledgments
Appendix A Properties of the Multinomial Logit Function
@@
Without loss of generality, we assume $\forall\mathcal{S}_{h,s,a},\exists s^{\prime}\in\mathcal{S}_{h,s,a}$ such that $\phi(s^{\prime}\>|\>s,a)=\mathbf{0}$. Otherwise, we can always define a new feature mapping $\phi^{\prime}(s^{\prime\prime}\>|\>s,a)=\phi(s^{\prime}\>|\>s,a)-\phi(s^{\prime\prime}\>|\>s,a)$ for any $s^{\prime\prime}\in\mathcal{S}_{h,s,a}$ such that $\phi^{\prime}(s^{\prime}\>|\>s,a)=\mathbf{0}$ and the transition induced by $\phi^{\prime}$ is the same as that induced by $\phi$. We denote this state as $\dot{s}_{h,s,a}$ and $\dot{\mathcal{S}}_{h,s,a}=\mathcal{S}_{h,s,a}\backslash\{\dot{s}_{h,s,a}\}$.
Recall the following definitions in the paper:
$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),$
$g_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$
$H_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.$
$\widehat{\Sigma}_{k,h}$ $=\frac{\kappa}{2}\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}+\lambda I_{d},$
$\mathcal{H}_{k,h}$ $=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda I_{d},$
@@
$H_{k,h}(\theta)$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
@@
Suppose $\lambda\geq 1$, define $\widetilde{\phi}_{k,h}^{s^{\prime}}=\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k+1,h})\phi_{k,h}^{s^{\prime\prime}}$, for any $k\in[K],h\in[H]$, the following statements hold:
$(\textnormal{I})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}^{2}\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)$
(II$)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)( I$
$(\textnormal{III})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
$(\textnormal{IV})\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
$(\textnormal{V})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
@@
Proof of statement (I).  By the definition that $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$, we have
$\det(\widehat{\Sigma}_{k+1,h})=\det(\widehat{\Sigma}_{k,h})\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{k,h}^{-1}}\Big{)}.$
Then, we get
$\sum_{i=1}^{k}\log\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{i,h}^{-1}}\Big{)}\leq\log\left(\frac{\det(\widehat{\Sigma}_{k+1,h})}{\det(\widehat{\Sigma}_{1,h})}\right)\leq d\log\left(1+\frac{kU}{\lambda d}\right),$ (12)
where the last inequality holds by determinant trace inequality in Lemma 16. Since $\lambda\geq 1$, we have $\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\leq 1$, thus, we have
$\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}$ $\leq\frac{2}{\kappa}\sum_{i=1}^{k}\min\bigg{\{}1,\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\bigg{\}}$
$\leq\frac{4}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{i,h}^{-1}}\Big{)}$
$\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right),$
where the first inequality holds by the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$ and the last inequality holds by (12).
Proof of statement (II).  By Lemma 5, we have $H_{k,h}(\theta)\succeq\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$. Thus, we have
$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k,h}+\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})p_{k,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{k+1,h})\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
Then, we get
$\det(\mathcal{H}_{i+1,h})\geq\det(\mathcal{H}_{i,h})\Big{(}1+\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}\Big{)}.$

### v3

@@ 6 Lower Bound
In this section, we establish the lower bound for MNL mixture MDPs by presenting a novel reduction, which connects MNL mixture MDPs and the logistic bandit problem.
Consider the following logistic bandit problem (Faury et al., 2020): at each round $t\in[T]$, the learner selects an action $x_{t}\in\mathcal{X}$ and receives a reward $r_{t}$ sampled from Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*}))^{-1}$, where $\theta^{*}\in\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$ is the unknown parameter. The learner aims to to minimize the regret: ${\operatorname*{Reg}}^{\mathtt{LogB}}(T)=\max_{x\in\mathcal{X}}\sum_{t=1}^{T}\mu(x^{\top}\theta^{*})-\sum_{t=1}^{T}\mu(x_{t}^{\top}\theta^{*})$.
Theorem 3.
@@
Corollary 1 also implies a problem-independent lower bound of $\Omega(dH\sqrt{K})$ directly. Corollary 1 can be proved by combining Theorem 3 and the $\Omega(d\sqrt{T\kappa^{*}})$ lower bound for logistic bandits with infinite arms by Abeille et al. (2021). To the best of our knowledge, a lower bound for logistic bandits with finite arms has not been established, which is beyond the scope of this work. This absence leaves the lower bound for MNL mixture MDPs with a finite action space open through this reduction. However, after the submission of our work to arXiv (Li et al., 2024a), a follow up work by Park et al. (2024) proposed a new reduction that bridges MNL mixture MDPs with linear mixture MDPs by approximating MNL functions to linear functions. Leveraging this new reduction, they established a problem-independent $\Omega(dH^{3/2}\sqrt{K})$ lower bound for the finite action setting. This achievement confirms that our result is optimal in $d$ and $K$, only loosing by an $\mathcal{O}(H^{1/2})$ factor.
Dependence on $H$.  By the discussion in Remark 5, we note that our result is optimal with respect to $d$ and $K$, but loosing by an $\mathcal{O}(H^{1/2})$ factor. We discuss the challenges in improving the dependence on $H$. Notably, MNL mixture MDPs can be viewed as a generalization of linear mixture MDPs (Ayoub et al., 2020; Zhou et al., 2021). The pioneering work by Ayoub et al. (2020) achieved a regret bound of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K})$ for linear mixture MDPs, which matches our results in Theorem 1, differing only on the lower-order term. Later, Zhou et al. (2021) enhanced the dependence on $H$ and attained an optimal regret bound of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$. This was made possible by recognizing that the value function in linear mixture MDPs is linear, allowing for direct learning of the value function while incorporating variance information. In contrast, the value function for MNL mixture MDPs does not conform to a specific structure, posing a significant challenge in using the variance information of value functions. Thus, it remains open whether similar improvements on $H$ are attainable for MNL mixture MDPs.
@@ 7 Conclusion and Future Work
In this work, we addressing both statistical and computational challenges for MNL mixture MDPs, which leverage MNL function approximation to ensure valid probability distributions. Specifically, we propose a statistically efficient algorithm that achieve a regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$, eliminating the dependence on $\kappa^{-1}$ in the dominant term for the first time. Then, we introduce a computationally enhanced algorithm that achieves the same regret but with only constant cost. Finally, we establish the first lower bound for this problem, justifying the optimality of our results in $d$ and $K$.
There are several interesting directions for future work. First, there still exists a gap between the upper and lower bounds. How to close this gap remains an open problem. Besides, we focuses on stationary rewards in this work, extending MNL mixture MDPs to the non-stationary settings and studying the dynamic regret (Wei and Luo, 2021; Zhao et al., 2022; Li et al., 2023) is also an important direction.
@@ Acknowledgments
Appendix A Notations
In this section, we collect the notations used in the paper in Table 2.
Table 2: Notations used in the regret analysis.
Notation Definition and description
$\ell_{k,h}(\theta)$ $\triangleq-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta)$, per-episode loss function at episode $k$ and stage $h$
$g_{k,h}(\theta)$ $\triangleq\nabla\ell_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, gradient of loss $\ell_{k,h}(\theta)$
$H_{k,h}(\theta)$ $\triangleq\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime},\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
$\mathcal{L}_{k,h}(\theta)$ $\triangleq\sum_{i=1}^{k-1}\ell_{i,h}(\theta)+\frac{\lambda_{k}}{2}\lVert\theta\rVert_{2}^{2}$, the cumulative MLE loss
$\mathcal{G}_{k,h}(\theta)$ $\triangleq\nabla\mathcal{L}_{k,h}(\theta)=\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta)-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}+\lambda_{k}\theta$, gradient of $\mathcal{L}_{k,h}(\theta)$
$\mathcal{H}_{k,h}(\theta)$ $\triangleq\nabla^{2}\mathcal{L}_{k,h}(\theta)=\sum_{i=1}^{k-1}H_{i,h}(\theta)+\lambda_{k}I_{d}$, Hessian of MLE loss $\mathcal{L}_{k,h}(\theta)$
$\widehat{\theta}_{k,h}$ $\triangleq\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\mathcal{L}_{k,h}(\theta)$, the MLE estimator at episode $k$ and stage $h$
$\mathcal{H}_{k,h}$ $\triangleq\mathcal{H}_{k,h}(\widetilde{\theta}_{i+1,h})=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda_{k}I_{d}$, the cumulative look ahead Hessian
~$\widetilde{\mathcal{H}}_{k,h}$ $\triangleq\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$, the sum of look ahead Hessian and the Hessian of current loss
$\widetilde{\theta}_{k+1,h}$ $\triangleq\operatorname*{arg\,min}_{\theta\in\Theta}\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}^{2}$, the OMD estimator
Appendix B Properties of Multinomial Logit Function
@@
Without loss of generality, we assume $\forall\mathcal{S}_{h,s,a},\exists\dot{s}_{h,s,a}\in\mathcal{S}_{h,s,a}$ such that $\phi(\dot{s}\>|\>s,a)=\mathbf{0}$. Otherwise, we can always define a new feature mapping $\phi^{\prime}(s^{\prime\prime}\>|\>s,a)=\phi(s^{\prime}\>|\>s,a)-\phi(s^{\prime\prime}\>|\>s,a)$ for any $s^{\prime\prime}\in\mathcal{S}_{h,s,a}$ such that $\phi^{\prime}(s^{\prime}\>|\>s,a)=\mathbf{0}$ and the transition kernel induced by $\phi^{\prime}$ is the same as that induced by $\phi$. Furthermore, We denote the set $\dot{\mathcal{S}}_{h,s,a}=\mathcal{S}_{h,s,a}\backslash\{\dot{s}_{h,s,a}\}$.
First, we introduce the definition of self-concordant-like functions and demonstrate that the MNL loss function is self-concordant-like.
Definition 3 (Self-concordant-like function, Tran-Dinh et al. [2015]).
A convex function $f\in\mathcal{C}^{3}\left(\mathbb{R}^{m}\right)$ is $M$-self-concordant-like function with constant $M$ if:
$\left|\psi^{\prime\prime\prime}(s)\right|\leqslant M\|\mathbf{b}\|_{2}\psi^{\prime\prime}(s).$
for $s\in\mathbb{R}$ and $M>0$, where $\psi(s):=f(\mathbf{a}+s\mathbf{b})$ for any $\mathbf{a},\mathbf{b}\in\mathbb{R}^{m}$.
Proposition 1.
The per-episode MNL loss $\ell_{k,h}(\theta)$ and the cumulative MNL loss $\mathcal{L}_{k,h}(\theta)$ are both $3\sqrt{2}$-self-concordant-like for all $k\in[K],h\in[H]$.
Proof.
By proposition B.1 in Lee and Oh [2024], the per-episode MNL loss function $\ell_{k,h}(\theta)$ is $3\sqrt{2}$-self-concordant-like. Then, the cumulative MNL loss function $\mathcal{L}_{k,h}(\theta)$ is the sum of self-concordant-like functions and a quadratic function, it is also $3\sqrt{2}$-self-concordant-like. ∎
Lemma 4 (Zhang and Sugiyama [2023, Lemma 1]).
Let $\ell(\mathbf{z},y)=\sum_{k=0}^{K}\mathbf{1}\{y=k\}\cdot\log\left(\frac{1}{[\sigma(\mathbf{z})]_{k}}\right)$ where $\sigma(\mathbf{z})_{k}=\frac{e^{z_{k}}}{\sum_{j=0}^{K}e^{z_{j}}}$, $\mathbf{a}\in[-C,C]^{K}$, $y\in\{0\}\cup[K]$ and $\mathbf{b}\in\mathbb{R}^{K}$ where $C>0$. Then, we have
$\ell(\mathbf{a},y)\geq\ell(\mathbf{b},y)+\nabla\ell(\mathbf{b},y)^{\top}(\mathbf{a}-\mathbf{b})+\frac{1}{\log(K+1)+2(C+1)}(\mathbf{a}-\mathbf{b})^{\top}\nabla^{2}\ell(\mathbf{b},y)(\mathbf{a}-\mathbf{b}).$
Then, we show the Hessian of the MNL loss function is positive semi-definite.
@@
$H_{k,h}(\theta)=$ $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
@@
Next, we show several concentration inequalities commonly used in the analysis.
@@
Suppose $\lambda_{k}\geq 1$, for any $k\in[K],h\in[H]$, for the quantities in Table 2 and define
$\bar{\phi}_{k,h}^{s^{\prime}}=\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\theta_{h}^{*})\phi_{k,h}^{s^{\prime\prime}},\quad\widetilde{\phi}_{k,h}^{s^{\prime}}=\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k+1,h})\phi_{k,h}^{s^{\prime\prime}}.$
Then, the following statements hold:
$(\textnormal{I})\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta_{h}^{*})\lVert\bar{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}(\theta_{h}^{*})}^{2}\leq 2d\log\left(1+\frac{k}{\lambda_{k}d}\right)$
(II$)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda_{k}d}\right)( I$
$(\textnormal{III})\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda_{k}d}\right)$
$(\textnormal{IV})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}(\theta)}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right),\forall\theta\in\Theta$
$(\textnormal{V})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right).$
@@
Proof of statement (I).  By the definition of $H_{k,h}(\theta)$, we have
$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}]-\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}]\big{(}\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime\prime}}]\big{)}^{\top}$
$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}\big{[}(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})^{\top}\big{]}$ (12)
Thus, we have $H_{i,h}(\theta_{h}^{*})\succeq\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})(\bar{\phi}_{i,h}^{s^{\prime}})(\bar{\phi}_{i,h}^{s^{\prime}})^{\top}$. Then, we get
$\mathcal{H}_{i+1,h}(\theta_{h}^{*})\succeq\mathcal{H}_{i,h}(\theta_{h}^{*})+\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})(\bar{\phi}_{i,h}^{s^{\prime}})(\bar{\phi}_{i,h}^{s^{\prime}})^{\top}$
As a result, we have
$\det(\mathcal{H}_{i+1,h}(\theta_{h}^{*}))\geq\det(\mathcal{H}_{i,h}(\theta_{h}^{*}))\Big{(}1+\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})\lVert\bar{\phi}_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}(\theta_{h}^{*})}\Big{)}.$